OpenStudy (anonymous):

How would I find the value of a function at a certain point given the intergral?

OpenStudy (anonymous):

I'll type the actual problem out here.

OpenStudy (anonymous):

\[F(x)=\int\limits_{0}^{x}dt \div(t^2+9)\]

OpenStudy (anonymous):

WHat would F(2) be?

OpenStudy (anonymous):

you need to solve for integral and then plug in x = 2. \[F'(t) = \frac{ 1 }{ t^2 + 9 }\] F(t) is the integral of F'(t) and plugging in 0 to x by FTC turn it into F(x) plug in 2 after ^_^

OpenStudy (anonymous):

Does it matter that in my problem it has F(x)= rather than F(t)=?

OpenStudy (anonymous):

not at all since it is the integral from 0 to x

OpenStudy (anonymous):

by FTC: \[F(x) = \int\limits_{0}^{x}f'(t) dt\]

OpenStudy (anonymous):

Oh! right, that makes total sense. Thank you!

OpenStudy (anonymous):

And If I wanted to find the derivative I just plug in x for t right?

OpenStudy (anonymous):

not sure the derivative of which function you mean

OpenStudy (anonymous):

the derivative of the intergral.

OpenStudy (anonymous):

F'(x) in this case.

OpenStudy (anonymous):

for some reason i'm not understanding the question. sorry ^_^